# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ 2.1 Controller’s Policy
Similar to classical MDPs, a history-dependent controller in the context of a robust MDP takes (randomized) actions based on the state-action sequence until the current state. Formally, the controller’s history $\left\{{\mathbf{H}_{t}:t\geq 0}\right\}$ is the $t$ indexed family of truncated sample paths
@@
We also define the random element $H_{t}:\Omega\rightarrow\mathbf{H}_{t}$ by point evaluation $H_{t}(\omega)=h_{t}$. Given a prescribed subset of probability measures $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$, a controller policy $\pi$ is a sequence of decision rules $\pi=(d_{0},d_{1},d_{2},\dots)$. Each decision rule $d_{t}$, indexed by $t$, is a measure valued function $d_{t}:\mathbf{H}_{t}\rightarrow\mathcal{Q}$. We call the set $\mathcal{Q}$ the action distributions of the controller.
@@
$\Pi^{\mathrm{C}}_{\mathrm{H}}:=\left\{{\pi=(d_{0},d_{1},\dots):d_{t}\in\mathbf{H}_{t}\rightarrow\mathcal{Q},\forall t\geq 0}\right\}.$ (2.1)
Note that the superscript “$\mathrm{C}$” stands for constrained. We also suppress the dependence on $\mathcal{Q}$ to simplify the notation.
A decision rule $d_{t}$ is Markov if for all $h_{t},h^{\prime}_{t}\in\mathbf{H}_{t}$ s.t. $s_{t}=s^{\prime}_{t}$, then $d_{t}(h_{t})=d_{t}(h^{\prime}_{t})$. Therefore, a Markov decision rule $d_{t}$ is indifferent to the state-action sequence $(s_{0},a_{0},\dots,s_{t-1},a_{t-1})$. Consequently, we can express $d_{t}(s_{t})=d_{t}(h_{t})$ when $d_{t}$ is Markov, employing an abuse of notation for the sake of simplicity. Then, define the set of $\mathcal{Q}$-constrained Markov controller policies is defined as
$\Pi^{\mathrm{C}}_{\mathrm{M}}:=\left\{{\pi=(d_{0},d_{1},\dots):d_{t}\in\mathbf{H}_{t}\rightarrow\mathcal{Q}\text{ is Markov},\forall t\geq 0}\right\}.$
A contoller’s policy $\pi=(d_{0},d_{1},\dots)$ is called Markov time-homogeneous (a.k.a. static as in Iyengar (2005)) if for any $t_{1},t_{2}\geq 0$ and $h_{t_{1}}\in\mathbf{H}_{t_{1}},h^{\prime}_{t_{2}}\in\mathbf{H}_{t_{2}}$ s.t. $s_{t_{1}}=s^{\prime}_{t_{2}}$, then $d_{t_{1}}(h_{t_{1}})=d_{t_{2}}(h^{\prime}_{t_{2}})$. In the case of a Markov time-homogeneous policy, the decision rule remains consistent regardless of the time and history once a state $s$ is specified—hence the term “static”. Consequently, such a policy can be succinctly represented as $\pi=(d,d,d,\ldots)$, where $d:S\rightarrow\mathcal{P}(\mathcal{A})$. Then, the set of $\mathcal{Q}$-constrained Markov time-homogeneous (or static) controller policies is defined as
$\Pi^{\mathrm{C}}_{\mathrm{S}}:=\left\{{\pi=(d,d,\dots):d\in S\rightarrow\mathcal{Q}}\right\}.$
Of particular interests in RL applications are the following two special type of constraints for controller’s decision rule. First, when $\mathcal{Q}=\mathcal{P}(\mathcal{A})$ the set of probability measures on $(A,\mathcal{A})$, then we have the fully randomized (or unconstrained) history-dependent/Markov/Markov time-homogeneous policy classes: $\Pi_{\mathrm{H}}/\Pi_{\mathrm{M}}/\Pi_{\mathrm{S}}$. Second, if $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}:=\left\{{\delta_{\left\{{a}\right\}}:a\in A}\right\}$, then we denote the deterministic history-dependent/Markov/Markov time-homogeneous policy classes as $\Pi^{\mathrm{D}}_{\mathrm{H}}/\Pi^{\mathrm{D}}_{\mathrm{M}}/\Pi^{\mathrm{D}}_{\mathrm{S}}$ respectively.
@@ 2.2 Adversary’s Policy
Similar to the controller, an adversarial policy $\kappa:=(\kappa_{0},\kappa_{1},\kappa_{2},\dots)$ is also a sequence of measure-valued functions of the history $\kappa_{t}:\mathbf{G}_{t}\rightarrow\mathcal{P}(\mathcal{S})$. For any $t\geq 0$, we call $\kappa_{t}$ the adversarial decision rule at $t$. Under this formulation, the largest set of adversarial policies $\mathrm{K}_{\mathrm{H}}$, a.k.a. the set of unconstrained history-dependent adversarial policies is
$\mathrm{K}_{\mathrm{H}}:=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathbf{G}_{t}\rightarrow\mathcal{P}(\mathcal{S}),\forall t\geq 0}\right\}.$
Analogue to the controller-side definitions, we call an adversarial decision rule $\kappa_{t}$ Markov if $\kappa_{t}$ if for all $g_{t},g^{\prime}_{t}\in\mathbf{G}_{t}$ s.t. $(s_{t},a_{t})=(s^{\prime}_{t},a_{t}^{\prime})$, then $\kappa_{t}(g_{t})=\kappa_{t}(g^{\prime}_{t})$. So, a Markov adversarial decision rule $\kappa_{t}$ is indifferent to $g_{t-1}=(s_{0},\dots,a_{t-1})$ given $(s_{t},a_{t})$. Thus, by an abuse of notation, we can write
$\kappa_{t}(s_{t},a_{t})=\kappa_{t}(g_{t}).$ (2.2)
Moreover, an adversary’s policy $\kappa=(\kappa_{0},\kappa_{1},\dots)$ is called Markov time-homogeneous (or static) if for all $t_{1},t_{2}\geq 0$ and $g_{t_{1}}\in\mathbf{G}_{t_{1}},g^{\prime}_{t_{2}}\in\mathbf{G}_{t_{2}}$ s.t. $(s_{t_{1}},a_{t_{1}})=(s^{\prime}_{t_{2}},a^{\prime}_{t_{2}})$, then $\kappa_{t_{1}}(g_{t_{1}})=\kappa_{t_{2}}(g^{\prime}_{t_{2}})$. So, for a Markov time-homogeneous policy, once a state action pair $(s,a)$ is given, the distribution of the next state is the same regardless of the time and history—hence the term “static”. Again, such policies can be represented as $(\kappa,\kappa,\kappa,\dots)$ where $\kappa:S\times A\rightarrow\mathcal{P}(\mathcal{S})$.
In environments where a robust MDP serves as a viable model, it becomes necessary to limit the adversary’s power to a relatively constrained set of policies. This restriction is crucial because an adversary with an unconstrained history-dependent policy class tends to be overly potent, rendering the controller’s actions ineffective. Therefore, allowing unrestricted adversarial policy choices in the robust MDP model can result in excessively conservative controller’s policies and overly pessimistic estimates of value. On the other hand, in order to achieve reinforcement learning and optimal control, it is desirable for the optimal value to satisfy a dynamic programming principle (DPP), a.k.a. Bellman equation. We will provide a rigorous definition of this principle in the forthcoming sections. These two factors, i.e., restricting the adversary’s power and adhering to the Bellman equation, constitute the primary motivation behind the concept of “rectangularity” as introduced in Iyengar (2005). This notion has subsequently been extended and generalized in works such as Wiesemann et al. (2013) and Goyal and Grand-Clément (2023).
We first introduce the notion of rectangularity considered by González-Trejo et al. (2002), Iyengar (2005), and Nilim and El Ghaoui (2005). In contemporary terminology, this concept is referred to as SA-rectangularity.
@@ 2.2.1 SA-Rectangular Set of Adversary’s Policies
For each $(s,a)\in S\times A$, let $\mathcal{P}_{s,a}\subset{\mathcal{P}(\mathcal{S})}$ be a prescribed subset of probability measure, which can be equivalently seen as a subset of row vectors in $\mathbb{R}^{|S|}$. We first consider the SA-rectangular marginal set of history-dependent adversarial decision rules
$\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(t):=\left\{{\kappa_{t}:\kappa_{t}(g_{t-1},s,a)\in\mathcal{P}_{s,a},\forall(s,a)\in S\times A,g_{t-1}\in\mathbf{G}_{t-1}}\right\}.$ (2.3)
$\mathrm{K}^{\mathrm{SA}}_{\mathrm{M}}(t)$ and $\mathrm{K}^{\mathrm{SA}}_{\mathrm{S}}(t)$ can be analogously defined when $\kappa_{t}$ being Markov or Markov time-homogeneous. To avoid introducing unused notations, we omit the details. The SA-rectangular set of history-dependent adversarial policies is
$\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}:=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(t),\forall t\geq 0}\right\}=\prod_{t\geq 0}\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(t).$
@@
$\mathrm{K}_{\mathrm{M}}^{\mathrm{SA}}$ :=$:=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(t)\text{ is Markov},\forall t\geq 0}\right\}$ (2.4)
$\mathrm{K}_{\mathrm{S}}^{\mathrm{SA}}$ :=$:=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(t)\text{ is Markov time-homogeneous},\forall t\geq 0}\right\}.$
@@
$\mathrm{K}_{\mathrm{M}}^{\mathrm{SA}}$ =$=\left\{{(\kappa_{0},\kappa_{1},\dots):\kappa_{t}=\left\{{\kappa_{t}(s,a):(s,a)\in S\times A}\right\},\kappa_{t}(s,a)\in\mathcal{P}_{s,a},\forall(s,a)\in S\times A,t\geq 0}\right\},$ (2.5)
$\mathrm{K}_{\mathrm{S}}^{\mathrm{SA}}$ =$=\left\{{(\kappa,\kappa,\dots):\kappa=\left\{{\kappa(s,a):(s,a)\in S\times A}\right\},\kappa(s,a)\in\mathcal{P}_{s,a},\forall(s,a)\in S\times A}\right\}.$
We note that to preserve the mathematical rigor in later constructions (e.g. definition (2.15)), (2.5) should be understood in terms of (2.4). A benefit of the representation in (2.5) is that it has a more natural interpretation and hence widely used in the literature; e.g. (González-Trejo et al., 2002; Iyengar, 2005; Nilim and El Ghaoui, 2005), and formulation (2.21) in Li and Shapiro (2023). Take the Markov setting as an example, (2.5) can be interpreted as the adversary choosing transition kernel
$\kappa_{t}=\left\{{\kappa_{t}(s,a):(s,a)\in S\times A}\right\}\in\prod_{(s,a)\in S\times A}\mathcal{P}_{s,a}=:\mathcal{P}^{\mathrm{SA}}$
for each instance of time. The product set $\mathcal{P}^{\mathrm{SA}}$ is referred to as a SA-rectangular ambiguity set of transition kernels. In this paper, we will take on the notation in (2.4) as it expresses history-dependence in an elegant way. We emphasize that this digression is only meant for a comparison with the terminologies and formulations in the literature.
@@ 2.2.2 S-Rectangular Set of Adversary’s Policies
We also consider the notion of S-rectangularity. To our knowledge, this concept is introduced to the robust MDP literature independently in Le Tallec (2007) and Wiesemann et al. (2013) as a way to further limit the power of the adversary while preserving the dynamic programming principle. As exemplified in Example 1, an S-rectangular adversary could be a more natural assumption in many relevant robust MDP models in the space of operations research and management sciences in comparison to SA-rectangularity. This is often due to that an SA-rectangular adversary has too much freedom, and thus has the potential to violate some natural constraint of the modeled environment.
Instead of freely choosing $\kappa_{t}(g_{t-1},s,a)\in\mathcal{P}_{s,a}$, in the S-rectangular setting, for a fixed $s\in S$, choosing adversarial decision rule $\kappa_{t}(g_{t-1},s,a)=p_{s,a}$ for one $a\in A$ could affect the possible choice of $\kappa_{t}(g_{t-1},s,a^{\prime})$ for any other $a^{\prime}\in A$. Concretely, fix a prescribed set of measure-valued functions $\mathcal{P}_{s}\subset\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$ for every $s\in S$. An element $p_{s}\in\mathcal{P}_{s}$ can be equivalently seen as a $\mathbb{R}^{|A|\times|S|}$ matrix, where each row is a probability vector on $S$. Define the S-rectangular marginal set of history-dependent adversarial decision rules as
$\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(t):=\left\{{\kappa_{t}:\kappa_{t}(g_{t-1},s,\cdot)\in\mathcal{P}_{s},\forall s\in S,g_{t-1}\in\mathbf{G}_{t-1}}\right\}$ (2.6)
and the S-rectangular set of history-dependent/Markov/Markov time-homogeneous adversarial policies
$\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ :=$:=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(t),\forall t\geq 0}\right\}$ (2.7)
$\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$ :=$:=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(t)\text{ is Markov},\forall t\geq 0}\right\}$
$\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ :=$:=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(t)\text{ is Markov time-homogeneous},\forall t\geq 0}\right\}.$
@@
$\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$ =$=\left\{{(\kappa_{0},\kappa_{1},\dots):\kappa_{t}=\left\{{\kappa_{t}(s,a):(s,a)\in S\times A}\right\},\kappa_{t}(s,\cdot)\in\mathcal{P}_{s},\forall s\in S,t\geq 0}\right\},$
$\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ =$=\left\{{(\kappa,\kappa,\dots):\kappa=\left\{{\kappa(s,a):(s,a)\in S\times A}\right\},\kappa(s,\cdot)\in\mathcal{P}_{s},\forall s\in S}\right\}.$
Again, this should be understood as a re-interpretation of (2.7). In this representation, the adversary chooses transition kernels
$\kappa_{t}\in\prod_{s\in S}\mathcal{P}_{s}=:\mathcal{P}^{\mathrm{S}}$
instead of history dependent-decision rules, where $\mathcal{P}^{\mathrm{S}}$ is known as a S-rectangular ambiguity set.
With these definitions we are now ready to introduce a modeling environment where the S-rectangular set of adversarial decision rules can lead to a more natural implementation. Here, we also exemplify a S-rectangular ambiguity set that is not SA-rectangular.

### v4

@@ 2.1 Controller’s Policy
Similar to classical MDPs, a history-dependent controller in the context of an RMDP takes (randomized) actions based on the state-action sequence until the current state. Formally, the controller’s history $\left\{{\mathbf{H}_{t}:t\geq 0}\right\}$ is the $t$ indexed family of truncated sample paths
@@
We also define the random element $H_{t}:\Omega\rightarrow\mathbf{H}_{t}$ by point evaluation $H_{t}(\omega)=h_{t}$.
Given a prescribed subset of probability measures $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$, a controller policy $\pi$ is a sequence of decision rules $\pi=(\pi_{0},\pi_{1},\pi_{2},\dots)$. Each decision rule $\pi_{t}$, indexed by $t$, is a measure valued function $\pi_{t}:\mathbf{H}_{t}\rightarrow\mathcal{Q}$ and is represented in the conditional distribution format as $\pi_{t}(a|h_{t})\in[0,1]$ and $\sum_{a\in A}\pi_{t}(a|h_{t})=1$ . We call the set $\mathcal{Q}$ the action distributions of the controller.
@@
$\Pi_{\mathrm{H}}(\mathcal{Q}):=\left\{{\pi=(\pi_{0},\pi_{1},\dots):\pi_{t}\in\left\{{\mathbf{H}_{t}\rightarrow\mathcal{Q}}\right\},\forall t\geq 0}\right\}.$ (2.1)
A decision rule $\pi_{t}$ is Markov if for all $h_{t},h^{\prime}_{t}\in\mathbf{H}_{t}$ s.t. $s_{t}=s^{\prime}_{t}$, then $\pi_{t}(\cdot|h_{t})=\pi_{t}(\cdot|h^{\prime}_{t})$. Therefore, a Markov decision rule $\pi_{t}$ is indifferent to the state-action sequence $(s_{0},a_{0},\dots,s_{t-1},a_{t-1})$. Consequently, we can express $\pi_{t}(\cdot|s_{t})=\pi_{t}(\cdot|h_{t})$ when $\pi_{t}$ is Markov, employing an abuse of notation for the sake of simplicity. Then, define the set of $\mathcal{Q}$-constrained Markov controller policies is defined as
$\Pi_{\mathrm{M}}(\mathcal{Q}):=\left\{{\pi=(\pi_{0},\pi_{1},\dots):\pi_{t}\in\left\{{\mathbf{H}_{t}\rightarrow\mathcal{Q}}\right\}\text{ is Markov},\forall t\geq 0}\right\}.$
A contoller’s policy $\pi=(\pi_{0},\pi_{1},\dots)$ is called Markov time-homogeneous (a.k.a. static as in Iyengar, [28]) if for any $t_{1},t_{2}\geq 0$ and $h_{t_{1}}\in\mathbf{H}_{t_{1}},h^{\prime}_{t_{2}}\in\mathbf{H}_{t_{2}}$ s.t. $s_{t_{1}}=s^{\prime}_{t_{2}}$, then $\pi_{t_{1}}(\cdot|h_{t_{1}})=\pi_{t_{2}}(\cdot|h^{\prime}_{t_{2}})$. In the case of a Markov time-homogeneous policy, the decision rule remains consistent regardless of the time and history once a state $s$ is specified—hence the term “static”. Consequently, such a policy $\pi$ can be identified with a measure-valued function $\pi:S\rightarrow\mathcal{P}(\mathcal{A})$ by considering the policy as $(\pi,\pi,\pi,\ldots)$. Then, the set of $\mathcal{Q}$-constrained Markov time-homogeneous (or static) controller policies is defined as
$\Pi_{\mathrm{S}}(\mathcal{Q}):=\left\{{(\pi,\pi,\dots):\pi\in\left\{{S\rightarrow\mathcal{Q}}\right\}}\right\}.$
The underlying constraint set $\mathcal{Q}$ is usually clear from the context. Thus, we simplify the notation by writing $\Pi_{\mathrm{H}}\equiv\Pi_{\mathrm{H}}(\mathcal{Q})$, and the same convention is applied to $\Pi_{\mathrm{H}}$ and $\Pi_{\mathrm{S}}$.
Of particular interests to RL applications are the following two special type of constraints for controller’s decision rule. First, when $\mathcal{Q}=\mathcal{P}(\mathcal{A})$, $\Pi_{\mathrm{H}}(\mathcal{P}(\mathcal{A}))/\Pi_{\mathrm{M}}(\mathcal{P}(\mathcal{A}))/\Pi_{\mathrm{S}}(\mathcal{P}(\mathcal{A}))$ represent the fully randomized (or unconstrained) history-dependent/Markov/Markov time-homogeneous policy classes. Second, if $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}:=\left\{{\delta_{\left\{{a}\right\}}:a\in A}\right\}$, then $\Pi_{\mathrm{H}}(\mathcal{Q}^{\mathrm{D}})/\Pi_{\mathrm{M}}(\mathcal{Q}^{\mathrm{D}})/\Pi_{\mathrm{S}}(\mathcal{Q}^{\mathrm{D}})$ are the deterministic history-dependent/Markov/Markov time-homogeneous policy classes, respectively.
@@ 2.2 Adversary’s Policy
Similar to the controller, an adversarial policy $\kappa:=(\kappa_{0},\kappa_{1},\kappa_{2},\dots)$ is also a sequence of measure-valued functions of the history $\kappa_{t}:\mathbf{G}_{t}\rightarrow\mathcal{P}(\mathcal{S})$ and represented as $\kappa_{t}(s|g_{t})$. For any $t\geq 0$, we call $\kappa_{t}$ the adversarial decision rule at $t$. Under this formulation, the largest set of adversarial policies $\mathrm{K}_{\mathrm{H}}$, a.k.a. the set of unconstrained history-dependent adversarial policies is $\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\left\{{\mathbf{G}_{t}\rightarrow\mathcal{P}(\mathcal{S})}\right\},\forall t\geq 0}\right\}.$
Analogue to the controller-side definitions, we call an adversarial decision rule $\kappa_{t}$ Markov if for all $g_{t},g^{\prime}_{t}\in\mathbf{G}_{t}$ s.t. $(s_{t},a_{t})=(s^{\prime}_{t},a_{t}^{\prime})$, then $\kappa_{t}(\cdot|g_{t})=\kappa_{t}(\cdot|g^{\prime}_{t})$. So, a Markov adversarial decision rule $\kappa_{t}$ is indifferent to $g_{t-1}=(s_{0},\dots,a_{t-1})$ given $(s_{t},a_{t})$. Thus, by an abuse of notation, we can write
$\kappa_{t}(\cdot|s_{t},a_{t})=\kappa_{t}(\cdot|g_{t}).$ (2.2)
Moreover, an adversary’s policy $\kappa=(\kappa_{0},\kappa_{1},\dots)$ is called Markov time-homogeneous (or static) if for all $t_{1},t_{2}\geq 0$ and $g_{t_{1}}\in\mathbf{G}_{t_{1}},g^{\prime}_{t_{2}}\in\mathbf{G}_{t_{2}}$ s.t. $(s_{t_{1}},a_{t_{1}})=(s^{\prime}_{t_{2}},a^{\prime}_{t_{2}})$, then $\kappa_{t_{1}}(\cdot|g_{t_{1}})=\kappa_{t_{2}}(\cdot|g^{\prime}_{t_{2}})$. So, for a Markov time-homogeneous policy, once a state action pair $(s,a)$ is given, the distribution of the next state is the same regardless of the time and history—hence the term “static”. Again, such policies can be represented as $(\kappa,\kappa,\kappa,\dots)$ where $\kappa:S\times A\rightarrow\mathcal{P}(\mathcal{S})$.
As motivated in the introduction, in environments where an RMDP serves as a viable model, it becomes necessary to restrict the adversary’s power to a constrained set of policies. On the other hand, a DPP, expressed through the Bellman equation (1.2), is crucial for certifying stationary Markov optimal policies and for enabling reinforcement learning and optimal control. These two considerations form the primary motivation behind the concept of rectangularity, as introduced in [28]. This notion has subsequently been extended and generalized in works such as [79] and [19].
We first introduce the notion of rectangularity considered by [18], [28], and [47]. In contemporary terminology, this concept is referred to as SA-rectangularity.
@@ 2.2.1 SA-Rectangular Set of Adversary’s Policies
[…]
$\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}},t):=\left\{{\kappa_{t}:\kappa_{t}(\cdot|g_{t-1},s,a)\in\mathcal{P}_{s,a},\forall g_{t-1}\in\mathbf{G}_{t-1},(s,a)\in S\times A}\right\}.$ (2.3)
$\mathrm{K}^{\mathrm{SA}}_{\mathrm{M}}(\mathcal{P}^{\mathrm{SA}},t)$ and $\mathrm{K}^{\mathrm{SA}}_{\mathrm{S}}(\mathcal{P}^{\mathrm{SA}},t)$ can be analogously defined when the adversary is restricted to be Markov or Markov time-homogeneous, respectively.
Then, the SA-rectangular set of history-dependent adversarial policies induced by $\mathcal{P}^{\mathrm{SA}}$ is
$\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}}):=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}},t),\forall t\geq 0}\right\}=\bigtimes_{t\geq 0}\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}},t).$
@@
$\mathrm{K}_{\mathrm{M}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}})$ $=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}},t)\text{ is Markov},\forall t\geq 0}\right\}$ (2.4)
$\mathrm{K}_{\mathrm{S}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}})$ $=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}},t)\text{ is Markov time-homogeneous},\forall t\geq 0}\right\}.$
Typically, the set $\mathcal{P}^{\mathrm{SA}}$ in the definition of adversarial policy classes is clear from the context. Therefore, we write $\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}\equiv\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}(\mathcal{P}^{\mathrm{SA}})$, with the same convention applied to $\mathrm{K}_{\mathrm{M}}^{\mathrm{SA}}$ and $\mathrm{K}_{\mathrm{S}}^{\mathrm{SA}}$ as well.
@@
$\mathrm{K}_{\mathrm{M}}^{\mathrm{SA}}$ $=\left\{{(\kappa_{0},\kappa_{1},\dots):\kappa_{t}=\left\{{\kappa_{t}(\cdot|s,a):(s,a)\in S\times A}\right\},\kappa_{t}(\cdot|s,a)\in\mathcal{P}_{s,a},\forall(s,a)\in S\times A,t\geq 0}\right\},$ (2.5)
$\mathrm{K}_{\mathrm{S}}^{\mathrm{SA}}$ $=\left\{{(\kappa,\kappa,\dots):\kappa=\left\{{\kappa(\cdot|s,a):(s,a)\in S\times A}\right\},\kappa(\cdot|s,a)\in\mathcal{P}_{s,a},\forall(s,a)\in S\times A}\right\}.$
We note that to preserve the consistency for later constructions (e.g. definition (2.13)), (2.5) should be understood in terms of (2.4). A benefit of the representation in (2.5) is that it has a more natural interpretation and hence widely used in the literature; e.g. [18, 28, 47], and formulation (2.21) in Li and Shapiro, [37]. Take the Markov setting as an example, (2.5) can be interpreted as the adversary choosing the transition kernel for each time $t$ as $p_{t}=\left\{{\kappa_{t}(\cdot|s,a)\in\mathcal{P}_{s,a}:(s,a)\in S\times A}\right\}\in\mathcal{P}^{\mathrm{SA}}.$ In this paper, we will take on the notation in (2.4) as it expresses history-dependence in an elegant way. We emphasize that this digression is only meant for a comparison with the terminologies and formulations in the literature.
@@ 2.2.2 S-Rectangular Set of Adversary’s Policies
We also consider the notion of S-rectangularity. To our knowledge, this concept is introduced to the RMDP literature independently in Le Tallec, [34] and Wiesemann et al., [79] as a way to further limit the power of the adversary while preserving the dynamic programming principle. As exemplified in Example 1, an S-rectangular adversary could be a more natural assumption in many relevant RMDP models in the space of operations research and management sciences in comparison to SA-rectangularity. This is often due to that an SA-rectangular adversary has too much freedom, and thus has the potential to violate some natural constraint of the modeled environment (e.g. see Example 1).
Instead of freely choosing $\kappa_{t}(\cdot|g_{t-1},s,a)\in\mathcal{P}_{s,a}$, in the S-rectangular setting, for a fixed $s\in S$, choosing adversarial decision rule $\kappa_{t}(\cdot|g_{t-1},s,a)=p_{s,a}$ for one $a\in A$ could affect the possible choice of $\kappa_{t}(\cdot|g_{t-1},s,a^{\prime})$ for any other $a^{\prime}\in A$. Concretely, fix a prescribed non-empty set of measure-valued functions $\mathcal{P}_{s}\subset\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$ for every $s\in S$. An element $p_{s}\in\mathcal{P}_{s}$ can be equivalently seen as a $\mathbb{R}^{|A|\times|S|}$ matrix, where each row is a probability vector on $S$. Also, we denote the collection of S-rectangular ambiguity sets as $\mathcal{P}^{\mathrm{S}}:=\bigtimes_{s\in S}\mathcal{P}_{s}$.
Define the S-rectangular marginal set of history-dependent adversarial decision rules induced by $\mathcal{P}_{s}$ as
$\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(\mathcal{P}^{\mathrm{S}},t):=\left\{{\kappa_{t}:\kappa_{t}(\cdot|g_{t-1},s,\cdot)\in\mathcal{P}_{s},\forall g_{t-1}\in\mathbf{G}_{t-1},s\in S}\right\}$ (2.6)
and the S-rectangular set of history-dependent/Markov/Markov time-homogeneous adversarial policies induced by $\mathcal{P}^{\mathrm{S}}$ as
$\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(\mathcal{P}^{\mathrm{S}})$ $=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(\mathcal{P}^{\mathrm{S}},t),\forall t\geq 0}\right\}$ (2.7)
$\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}(\mathcal{P}^{\mathrm{S}})$ $=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(\mathcal{P}^{\mathrm{S}},t)\text{ is Markov},\forall t\geq 0}\right\}$
$\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}(\mathcal{P}^{\mathrm{S}})$ $=\left\{{\kappa=(\kappa_{0},\kappa_{1},\dots):\kappa_{t}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}(\mathcal{P}^{\mathrm{S}},t)\text{ is Markov time-homogeneous},\forall t\geq 0}\right\}.$
As before, we will omit the dependence on $\mathcal{P}^{\mathrm{S}}$ when it is clear from the context.
@@
$\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$ $=\left\{{(\kappa_{0},\kappa_{1},\dots):\kappa_{t}=\left\{{\kappa_{t}(\cdot|s,a):(s,a)\in S\times A}\right\},\kappa_{t}(\cdot|s,\cdot)\in\mathcal{P}_{s},\forall s\in S,t\geq 0}\right\},$
$\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ $=\left\{{(\kappa,\kappa,\dots):\kappa=\left\{{\kappa(\cdot|s,a):(s,a)\in S\times A}\right\},\kappa(\cdot|s,\cdot)\in\mathcal{P}_{s},\forall s\in S}\right\}.$
Again, this should be understood as a re-interpretation of (2.7). In this representation, the adversary chooses transition kernels $p_{t}\in\mathcal{P}^{\mathrm{S}}$.